1. Suppose that a consumption equation is of the form.
C= 145 + 0.67 Yd where Yd is disposable income.

a. Interpret what the figures 145 and 0.67 imply in the consumption equation.

b. What is the associated saving functions? What do the intercept and slope mean?

c. Sketch the savings and the associated consumption function.

d. Compute the level of consumption if income Yd is GHS 200.
1
Expert's answer
2021-10-18T11:28:27-0400

Solution:

a.).Figure 145 = Autonomous consumption. This is the consumption expenditure that occurs when income levels are zero.

Figure 0.67 = Marginal Propensity to Consume (MPC). This is the proportion of additional income that an individual consumes.

 

b.). Saving function = Y – C

S = -C + (1 – b) Y

MPS = 1 – 0.67 = 0.33

Saving function = -145 + 0.33Y

The intercept refers to the autonomous consumption and the slope refers to the marginal propensity to consume and indicated induced consumption.

 

c.). The savings and the associated consumption graph is as follows:



 

d.). C = 145 + 0.67Yd

Yd = 200

C = 145 + 0.67 (200)

C = 145 + 134 = 279

The level of consumption = GHS 279